### v1

@@ 4.2 Limitations and Points of Vigilance
Beyond these technical and ethical concerns, using an LLM to generate agent behavior fundamentally alters the modeling process itself. Unlike traditional agent-based modeling, where the designer explicitly selects or constructs behavioral models based on theory or empirical evidence, LLM-driven agents bypass this stage: the behavioral model is implicit, embedded within billions of opaque parameters learned from large-scale data. As a result, the modeller relinquishes the ability to clearly define or justify the assumptions underlying agent behavior. This lack of epistemic control is especially problematic in scientific contexts, where the rationale for modeling choices must be transparent and subject to scrutiny. While LLMs may, in theory, reflect statistical patterns closer to real-world data, this advantage is undermined if the underlying mechanisms remain inaccessible and unvalidated.
4.3 LLM Applications: Interactive Use vs. Social Simulation Challenges
In contexts such as educational games, training simulations, or immersive environments (Taillandier et al. 2019b), LLMs offer direct and significant benefits. Here, the primary goal is often to create a ”believable” and ”engaging” experience rather than a perfectly accurate, verifiable replication of reality (Costa 2025). LLMs excel at simulating dynamic, personalized interactions with artificial agents, thereby significantly enhancing user engagement. They can provide adaptive learning pathways, tailoring content and activities to individual needs and preferences. This fosters critical thinking by presenting users with diverse scenarios and responses that encourage problem-solving and decision-making (Hu et al. 2025; Costa 2025). For instance, in a model on teacher-student interactions,LLMs can provide personalized feedback to learners, offering constructive criticism and guidance that is difficult to scale in traditional educational settings. They also contribute to the creation of immersive learning environments through gamification, where learning objectives are interwoven with interactive narratives and challenges (Hu et al. 2025; Costa 2025). The integration of LLMs with virtual and augmented reality holds promise for creating highly immersive and interactive learning experiences, allowing users to step into simulated environments and engage with AI agents as main characters (Costa 2025). Furthermore, LLMs can facilitate personalized learning paths, adapting content and activities to individual needs and preferences, and fostering human-AI collaboration where AI assists teachers with repetitive tasks, freeing them to focus on student-centric activities (Costa 2025).
However, for “pure social simulation” aimed at understanding or predicting complex societal phenomena, the application of LLMs is more complicated due to existing limitations (Chen et al. 2024b). The inherited biases from training data can lead to inaccurate or discriminatory simulation outcomes (Santurkar 2025). The black-box nature makes it difficult to interpret the underlying mechanisms of simulated social phenomena, hindering scientific explanation and theory building (Li 2024). The heavy computational cost limits the use of classical sensitivity analysis approaches, which are crucial for understanding model robustness and parameter influence in traditional social simulations (Wu et al. 2025). The absence of genuine inner psychology, emotions, and intent in LLMs means their simulated behaviors might not accurately reflect the complex decision-making processes of humans in real social contexts (Wang et al. 2025). The ”average persona” problem and lack of behavioral diversity limit the ability to simulate heterogeneous populations and emergent phenomena that arise from individual differences (Wu et al. 2025). Finally, ensuring alignment with real-world patterns, consistency over time, and reproducibility remains a significant challenge, making it difficult to establish the credibility of findings for scientific discovery (Wu et al. 2025). These challenges necessitate careful consideration of the research question and the appropriate level of fidelity required when deploying LLMs for pure social science inquiry.
4.4 Research Fronts and Future Improvements
Future research must prioritize addressing diversity, bias, sycophancy, alienness, and generalization issues in LLM-based simulations (Wu et al. 2025). To enhance realism, developing more diverse and personalized training datasets, including first-person narratives, psychological insights, and cultural experiences, is crucial. Incorporating external motivation structures could enable more human-like decision-making driven by goals, fears, or aspirations, moving beyond mere statistical prediction. Building richer virtual worlds where AI agents compete, collaborate, or survive in dynamic, resource-driven environments could lead to more lifelike decisions and emergent behaviors. New evaluation frameworks are needed to benchmark AI-driven social simulations more systematically. This includes, but is not limited to, metrics assessing psychological alignment with real human behaviors, in order to better calibrate and refine these models for complex societal dynamics. Simulations may be more valuable when focusing on collective patterns rather than individual trajectories, especially given current limitations in individual behavioral diversity (Wu et al. 2025).
Smaller language models (SLMs) are emerging as potentially more suitable and economical for many agentic systems, offering lower latency, reduced memory and computational requirements, and significantly lower operational costs while maintaining adequate task performance in constrained domains (Belcak et al. 2025). This shift could promote more responsible and sustainable AI deployment, addressing some of the environmental and economic concerns associated with large-scale LLMs (Belcak et al. 2025). Integrating LLMs with established theoretical perspectives from psychology, sociology, and decision science, as well as traditional agent-based models (ABMs), offers a powerful toolset for more nuanced and realistic models (Chen et al. 2024b). This hybrid approach allows researchers to combine the generative flexibility of LLMs with the structured analysis capabilities of ABMs. Hybridization can also take the form of architectural compromises — for instance, assigning a single LLM to generate responses for a group of agents, while injecting agent-specific noise or contextual variations to simulate individual behavioral diversity within the simulation. This strategy offers a pragmatic trade-off between computational efficiency and behavioral differentiation, especially in large-scale systems. Continued focus on ethical alignment is critical, especially as LLMs take on complex agentic roles involving decision-making with humans, ensuring that their behaviors align with human values and societal norms (Liu et al. 2024).
@@ 5 Conclusion
LLMs have demonstrated impressive capabilities in mimicking human-like conversation and passing complex psychological tests, including the Turing Test and certain Theory of Mind tasks (Jones & Bergen 2025; Kosinski 2023). These capabilities have spurred the development of advanced LLM-driven social simulation frameworks like Generative Agents (Smallville) and AgentSociety (Park 2023; Piao et al. 2025). However, these successes are rooted in sophisticated pattern recognition rather than genuine understanding, leading to significant limitations. These include inherent biases, a black-box nature, high computational costs, a lack of true inner psychology or behavioral diversity, and issues with hallucination and inconsistency (Schramowski 2024; Poeppel & Marantz 2025; Wu et al. 2025; Li 2024). While LLMs show direct utility and promise in interactive simulations and serious games for engagement and personalized learning (Costa 2025; Hu et al. 2025), their application in pure social simulation for accurate prediction or replication of complex human societies remains challenging and requires careful validation and boundary definition (Chen et al. 2024b; Zhou 2024).
The current landscape suggests a fundamental shift in the approach to social simulation. The limitations of pure LLM-based simulations—such as their inherent biases, black-box nature, computational demands, and tendency towards behavioral homogeneity (Wu et al. 2025; Li 2024)—underscore the enduring value of traditional Agent-Based Models (ABMs) for structured environments and rigorous analysis. This has led to a growing interest in integrating LLMs into classical simulation platforms like GAMA (Taillandier et al. 2019a) and NetLogo (Jimenez-Romero et al. 2025; Zhang et al. 2025a; Liu et al. 2025). This approach, often referred to as LLM-augmented ABM, offers a powerful toolset for researchers, allowing for more nuanced, realistic, and comprehensive models of complex systems and human behaviors (Chen et al. 2024b). Such mixed models can strategically leverage the strengths of LLMs, such as their natural language processing capabilities, context-aware processing, and adaptive behaviors, to enrich agent intelligence (Piao et al. 2025). Concurrently, traditional ABM frameworks provide the structured environment, explicit interaction rules, and mechanisms for rigorous analysis and validation (Piao et al. 2025; Wu et al. 2025; Zhang et al. 2025a). Examples of such integration are already emerging, with toolchains integrating LLMs with NetLogo via Python extensions for prompt-driven behavior generation (Jimenez-Romero et al. 2025). A similar integration effort is also underway in the GAMA platform (Nghi et al. 2025), which supports LLM-based behavior scripting through external APIs and Python bridges, allowing modelers to delegate selected decision-making components to language models while preserving the structure and analytical transparency of rule-based modeling. This integration also offers access to a wide set of models, enabling flexible experimentation and comparative evaluation of different LLM architectures within the same modeling environment. This hybrid approach allows modelers familiar with traditional platforms to harness LLM capabilities, fostering the development of models that combine the strengths of both paradigms.
Invoking large language models (LLMs) during agent-based simulations raises significant computational challenges, especially when such calls occur frequently or at scale, potentially limiting their practical usability in large experiments or real-time applications. In response to these constraints, the discussion about smaller language models (SLMs) as the ”future of agentic AI” for efficiency (Belcak et al. 2025) suggests a move towards modularity, where different models (LLM, SLM, rule-based) are chosen for specific sub-tasks based on their strengths and weaknesses. This indicates a paradigm shift in computational social science: instead of LLMs replacing traditional ABMs, they are becoming powerful components within a broader, more sophisticated modeling ecosystem. The future of social simulation will likely involve highly modular, hybrid architectures that strategically combine LLMs for complex, language driven behaviors with traditional rule-based or statistical models for structured interactions, large-scale dynamics, and rigorous analysis, potentially augmented by SLMs for efficiency in specific sub-tasks. This allows researchers to exploit the unique strengths of LLMs while mitigating their inherent limitations, leading to more robust, interpretable, and scientifically valuable social simulations. This also suggests a need for new methodologies to manage the complexity of these hybrid systems and validate their emergent behaviors.

### v2

@@ 4.2 The ”Black-Box” Problem and Interpretability
Beyond these ethical concerns, using an LLM to generate agent behavior fundamentally alters the scientific modeling process itself. Unlike traditional Agent-Based Modeling (ABM), where the designer explicitly constructs behavioral rules based on theory or empirical evidence, LLM-driven agents bypass this stage. The behavioral model becomes implicit, embedded within billions of opaque parameters learned from massive datasets. As a result, the modeller relinquishes the ability to clearly define or justify the assumptions underlying agent behavior. This lack of epistemic control is especially problematic in scientific contexts, where the rationale for modeling choices must be transparent and subject to scrutiny. While LLMs may, in theory, reflect statistical patterns closer to real-world data, this advantage is undermined if the underlying mechanisms remain inaccessible and unvalidated.
4.3 Systemic Bias and the ”Average Persona” Risk
While Section 2.3 highlighted the intrinsic biases of LLMs, their impact scales non-linearly when applied to multi-agent populations. The most critical risk for social simulation is systemic homogenization. As noted by [81], the statistical nature of LLMs drives them to converge toward a generic, safe, and ”WEIRD” behavioral mean. In a simulation context, this creates a ”Replicant Effect”: even when agents are assigned distinct demographic prompts, their actual decision-making logic collapses into a uniform mode, effectively erasing minority behaviors and extreme viewpoints. Moreover, this homogenization is paradoxically accompanied by ”Generative Exaggeration”. As demonstrated by [52], when LLMs are prompted to adopt specific personas to escape the ”average,” they often drift into caricature, amplifying stereotypical traits to the point of toxicity rather than maintaining behavioral nuance. This suggests that current models struggle to find a middle ground between the ”flatness” of the average persona and the ”parody” of specific roles.
This ”Average Persona” problem is exacerbated by the ”Data Scarcity Wall.” As discussed regarding the limitations of fine-tuning, simulating specific, under-represented populations—such as smallholder farmers in the Global South—is often impossible because the massive, digitized behavioral logs required for training simply do not exist. Consequently, relying on generic LLMs to simulate these groups risks replacing missing data with hallucinated stereotypes.
To mitigate this, future research must move beyond mere prompt engineering. Developing more diverse and personalized training datasets is crucial, specifically incorporating first-person narratives, oral histories, and qualitative psychological insights to capture the ”long tail” of human behavior. Furthermore, incorporating external motivation structures (e.g., BDI architectures) can help force agents to deviate from the statistical mean by driving decisions through explicit goals and fears rather than mere text prediction. Without these structural interventions, simulations will remain valuable for exploring collective patterns of the ”average” human [81], but will fail to capture the diversity essential for understanding complex societal dynamics.
4.4 Computational Constraints and Scalability
While LLMs offer cost-efficiency compared to human experiments, they introduce a drastic computational bottleneck compared to traditional Agent-Based Models (ABMs). The resource intensity of LLMs—both in fine-tuning and real-time inference—limits practical scalability [16]. Crucially, this high cost is not merely a logistical hurdle but a methodological barrier: it hinders the application of classical sensitivity analyses, which require running simulations thousands of times to assess parameter influence and model robustness [56, 81]. Unlike lightweight rule-based agents that permit exhaustive parameter sweeps on standard hardware, the computational latency of LLMs currently forces researchers to trade statistical rigor for behavioral complexity, often preventing the extensive replications necessary to validate stochastic models.
4.5 Ethics in Application: Serious Games vs. Predictive Science
Given the methodological barriers outlined above, it is crucial to delineate the appropriate domains of application for LLM-based agents.
4.5.1 Interactive Environments and Serious Games
In contexts such as educational games and training simulations [71], LLMs offer direct and significant benefits. Here, the primary goal is often to create a ”believable” and ”engaging” experience rather than a perfectly verifiable replication of social reality [16]. LLMs excel at simulating dynamic, personalized interactions, enhancing user engagement through adaptive learning pathways and gamified narratives [30].
However, as highlighted by recent critiques, this application requires a shift from cautious optimism to critical vigilance. The primary risk lies in ’negative training transfer’: unlike traditional rule-based simulators where every outcome is verified, an LLM-based agent can hallucinate factual errors with high confidence. Recent studies highlight that this ’authoritative misinformation’ can lead to superficial learning and cognitive over-reliance [88]. Furthermore, instructional quality is threatened by inherent biases; as demonstrated by [76], models often perpetuate harmful stereotypes in personalized content. Consequently, the deployment of LLMs in training necessitates rigorous ’red-teaming’ and grounded architectures (RAG) to ensure the simulation remains a tool for learning, not confusion.
4.5.2 From Prediction to Exploratory Modeling
For ”pure social simulation” aimed at scientific inquiry, the implications of the ”validity gap” and ”epistemic opacity” (discussed in Sections 4.1 and 4.2) are decisive. Because of the risks of hallucination and the ”average persona” bias, we argue that LLMs are currently ill-suited for confirmatory empirical research or precise forecasting (e.g., predicting specific election outcomes or policy impacts).
Instead, their most scientifically robust application lies in Exploratory Modeling [14]. In this paradigm, LLMs serve not as oracles of truth, but as engines for hypothesis generation. They allow researchers to rapidly explore a vast space of possible social interactions and identify emergent patterns that—while not necessarily predictive of the real world—reveal the logical consequences of specific behavioral assumptions. As [81] suggests, these simulations should be viewed as ”reasoning pumps” to test collective dynamics under diverse narrative constraints, rather than as digital twins of reality.
5 Towards Hybrid Constitutional Architectures
The current landscape suggests a fundamental shift in the approach to social simulation. While LLMs offer compelling opportunities—enabling the rapid simulation of thousands of interactions and uncovering emergent behaviors [4, 56]—the limitations of ”off-the-shelf” models underscore the enduring value of traditional Agent-Based Models (ABMs).
5.1 Integrating LLMs into Classical Frameworks
The future of social simulation does not lie in ”off-the-world” AI agents, but in the seamless integration of LLMs within structured platforms such as GAMA [70] and NetLogo [34]. This approach, which we define as LLM-augmented ABM, addresses the fundamental problem of groundedness: while LLMs manipulate abstract concepts, the ABM provides the ”world model”—the physical, spatial, and resource-based constraints—that dictates the actual viability of AI-proposed actions.
This synergy allows researchers to move beyond purely narrative modeling. Recent work, such as Gao et al. (2025) on the S3 system, demonstrates how individual behavior alignment can be synchronized with complex social network structures, ensuring that LLM expressiveness does not sacrifice macroscopic coherence [26]. Simultaneously, frameworks like AgentTorch [15] open the door to massive scalability through GPU acceleration, enabling simulations of entire populations where the ABM manages the physics of interaction while the LLM simulates context-sensitive psychology.
For instance, in an epidemiological model like COMOKIT [27, 69], an LLM could interpret evolving social norms or disinformation, while the ABM maintains the structural rigor of disease spread. This hybrid necessity is even more acute in high-stakes environments. [82] recently demonstrated this in disaster evacuation simulations: while the ABM manages the physical constraints of crowd density and velocity to prevent unrealistic movements, the LLM-enhanced agents simulate the cognitive panic and decision-making under uncertainty. This ”Physics-Cognition” split ensures that the simulation remains physically valid while capturing the psychological volatility of human crowds.
5.2 Cognitive Alignment: Fine-Tuning and its Limits
To correct the instability of generic models, a promising path lies in behavioral alignment. Binz et al. (2025) have demonstrated that it is possible to transform an LLM into a cognitively aligned agent (the Centaur model) by training it on massive datasets derived from human psychological experiments. This process allows the AI to replicate human learning curves and decision-making biases correlated with human neural signals [10].
However, this data-driven path hits a critical ”Data Scarcity Wall” as soon as one moves beyond Western (WEIRD) contexts. As highlighted by [60], when representative training data is absent, LLMs tend to fall back on generalized stereotypes to fill the gaps. Consequently, in social simulation where identity is relational and cultural, modeling the decisions of smallholder farmers in Vietnam is impossible through simple fine-tuning: the necessary digitized behavioral logs do not exist, leaving the model to rely on hallucinated or westernized proxies.
Pursuing a strategy exclusively based on data risks reinforcing the ”Average Avatar Bias”: an AI that perfectly mimics a mean Western human but fails entirely to represent the specificities of marginalized or culturally distinct groups. To break this impasse, alignment must shift from being purely statistical (data-driven) to being structural (theory-driven), justifying the transition toward ”constitutional” architectures.
5.3 Proposed Paradigm: Hybrid Constitutional Architectures
We argue that cognitive alignment in social simulation cannot remain purely data-driven; it must be structurally anchored in psychological theory. We propose the adoption of Hybrid Constitutional Architectures, a paradigm shift where the LLM’s generative freedom is bounded by a theoretical ’scaffold.’
In this stratified cognitive process, we move beyond the monolithic agent:
  • Heuristic Stratum (SLMs): High-frequency, routine behaviors (e.g., daily navigation, basic consumption) are delegated to Small Language Models (SLMs). This ensures computational scalability [9] and allows for ’local alignment’—where smaller models are fine-tuned on specific ethnographic data that would be diluted in larger architectures.
  • Reasoning Stratum (LLM as Proposal Engine): Larger models are invoked only for complex, non-routine social reasoning. Crucially, the LLM does not execute the final decision; it acts as a ’Proposal Engine’, generating multiple plausible behavioral trajectories based on the current social context.
  • The Theoretical Constitution: Every proposal is passed through a validation filter grounded in established frameworks, such as the Theory of Planned Behavior (TPB) [2] or Bounded Rationality principles [65]. For instance, a proposal that contradicts an agent’s established ’Subjective Norms’ or ’Perceived Behavioral Control’ is rejected or penalized. The constitutional layer must itself be treated as a model component subject to audit, since poorly specified normative constraints may reduce harmful outputs while simultaneously suppressing legitimate social heterogeneity.
  • External Memory Modules (ABM State): To resolve the Temporal Instability inherent to finite context windows, the architecture offloads long-term state (biography, resources, social graph) to the classical ABM environment. Rather than relying on the LLM to ’remember’ the entire history, relevant data is retrieved dynamically via Retrieval-Augmented Generation (RAG). This separation ensures that agents maintain consistent identities over thousands of simulation steps [55, 54].
To translate this conceptual architecture into an explicit implementation protocol, we formalize the decision-making process as a stratified execution loop that strictly arbitrates between the ABM, SLMs, and LLMs. Let $E_{t}$ denote the global environment state at simulation step $t$, managed by the classical ABM (e.g., GAMA, NetLogo) acting as the absolute ground truth, and $S_{i,t}$ the internal state of agent $i$.
Arbitration is governed by a context-shift function, $\Delta(E_{t},S_{i,t})$, which continuously evaluates environmental and social stability. If $\Delta(E_{t},S_{i,t})\leq\tau$ (where $\tau$ is a predefined stability threshold), the environment is considered stable. The agent defaults to the computationally lightweight Heuristic Stratum, generating a routine action via a Small Language Model or classical heuristics: $a_{\text{heur}}=\text{SLM}(E_{t},S_{i,t})$.
However, if the ABM detects a novel event, a social anomaly, or a critical threshold breach (e.g., a sudden policy change or resource depletion), $\Delta(E_{t},S_{i,t})>\tau$. Decision-making is then escalated to the Reasoning Stratum. Here, the LLM acts as a proposal engine, generating an intended behavioral trajectory based on context and retrieved memory ($M_{i}$): $a_{\text{prop}}=\text{LLM}(E_{t},S_{i,t},M_{i})$.
Crucially, this generated intention is never executed directly. It must pass through the Theoretical Constitution, formalized as a validation filter $\Phi(a,E_{t},S_{i,t})\in\{0,1\}$. Grounded in psychological frameworks like the Theory of Planned Behavior (TPB), $\Phi$ evaluates both the physical possibility of the action within the ABM and its normative plausibility for the specific agent. The final action execution follows a strict fallback mechanism:
$a_{\text{final}}=\begin{cases}a_{\text{prop}}&\text{if }\Delta>\tau\text{ and }\Phi(a_{\text{prop}},E_{t},S_{i,t})=1\\ a_{\text{heur}}&\text{otherwise (fallback or routine)}\end{cases}$ (1)
Finally, the ABM gathers all validated actions ($\mathbf{A}$), resolves physical conflicts, and updates the environment state for the next step: $E_{t+1}=f_{\text{ABM}}(E_{t},\mathbf{A})$. This formal protocol (summarized in Algorithm 1) ensures that the generative power of LLMs is rigorously bounded by causal and computational constraints.
Algorithm 1 Hybrid Constitutional Execution Loop (Single Agent Step)
1:Current environment state $E_{t}$, Agent internal state $S_{i,t}$, Stability threshold $\tau$
2:Updated environment state $E_{t+1}$
[…]
5:if $\delta>\tau$ then $\triangleright$ Escalate to LLM Reasoning
6:  $M_{i}\leftarrow\text{RAG}(S_{i,t})$ $\triangleright$ Retrieve contextual memory
7:  $a_{\text{prop}}\leftarrow\text{LLM}(E_{t},S_{i,t},M_{i})$ $\triangleright$ Generate behavioral proposal
8:  if $\Phi(a_{\text{prop}},E_{t},S_{i,t})==1$ then $\triangleright$ Constitutional Validation
9:   $a_{\text{final}}\leftarrow a_{\text{prop}}$
10:  else
11:   $a_{\text{final}}\leftarrow a_{\text{heur}}$ $\triangleright$ Fallback due to constitutional rejection
12:  end if
13:else
14:  $a_{\text{final}}\leftarrow a_{\text{heur}}$ $\triangleright$ Default to routine behavior
15:end if
16:Execute $a_{\text{final}}$ via ABM engine
[…]
To measure the analytical improvement of this hybrid approach against both classical rule-based ABMs and pure LLM-driven simulations, we propose a tripartite evaluation pipeline. First, Counterfactual Robustness: unlike classical ABMs that fail outside pre-coded rules, and pure LLMs that may hallucinate physically impossible actions, the hybrid model must be tested against unseen socio-environmental shocks to ensure responses are both socially plausible and physically grounded. Second, Behavioral Diversity: since pure LLMs tend to collapse into an ’Average Persona,’ the hybrid architecture—particularly through locally fine-tuned SLMs—must be quantitatively evaluated on its ability to preserve statistical heterogeneity and minority behaviors. Finally, Computational Scalability: a key metric of success is the ability to run exhaustive sensitivity analyses (e.g., Monte Carlo simulations with 1000+ runs). While pure LLM models face prohibitive API latency and costs, the stratified delegation to SLMs must demonstrate a significant reduction in computational overhead, restoring the rigorous statistical validation expected in traditional social simulation.
By treating the LLM as a creative engine constrained by a formal psychological constitution, we solve the ’black-box’ problem of intent. Decision transparency is regained: the modeler can verify if a choice was made due to a social norm (rule-based) or a creative adaptation (LLM-based). This synergy moves the field from ’surface realism’ to causal plausibility, enabling simulations that remain robust even when empirical data is scarce. Rather than replacing classical agent-based modeling, Hybrid Constitutional Architectures reposition LLMs as bounded cognitive modules within a transparent causal framework, thereby reconciling expressive flexibility with epistemic accountability.
@@ 6 Conclusion
LLMs have demonstrated impressive capabilities in mimicking human-like conversation and passing complex psychological tests [35, 40]. Yet, as this paper has explored, these successes are often rooted in sophisticated pattern recognition rather than genuine understanding. This leads to the ”Fluency Fallacy”: a state where models maintain high linguistic coherence even when hallucinating social facts, creating a ”surface realism” that can mislead researchers into mistaking narrative plausibility for causal validity.
The transition from individual generative agents to robust social simulations requires moving beyond the Turing Test toward stricter criteria: Behavioral Diversity (to avoid mode collapse towards an average persona) and Counterfactual Robustness (the ability to react realistically to unseen shocks) [81, 44]. While LLMs show direct utility in interactive simulations and serious games for engagement [16], their application in pure social science demands a critical re-evaluation of epistemic control.
Ultimately, the future of social simulation does not lie in a choice between the data-driven flexibility of LLMs and the theory-driven rigor of ABM, but in their synergy. By coupling the narrative transparency of Generative AI with the structural rigor of Hybrid Constitutional Architectures, we can achieve simulations that are not only plausible but causally valid and intuitively explainable. This hybrid paradigm ensures that even in data-scarce environments, artificial societies can serve as meaningful laboratories for understanding the deep-seated nuances of human decision-making.
Conflict of Interest Statement
The authors declare that the research was conducted in the absence of any commercial or financial relationships that could be construed as a potential conflict of interest.
Author Contributions
P.T. conceived the original structure and wrote the first draft of the manuscript. J.D.Z., A.G., B.G., N.Q.H., and A.D. substantially contributed to the conceptualization, provided critical feedback, and improved the manuscript. The development of this article was a truly collective effort. All authors discussed the concepts, read, and approved the final manuscript.